[figure: 9.7 cm]
FMR IN THE (La_{0.7}Ca_{0.3})_{1−x}Mn_{1+x}O_{3} FILMS.

V.P. Dyakonov^{a,b}, V. Shapovalov^{b}, E. Zubov^{b}, P. Aleshkevych^{a}, K. Dyakonov^{c} and H. Szymczak^{a} aInstitute of Physics, 02-668 Warszawa, Poland bA.A. Galkin Physico-Technical Institute, Donetsk 340114, Ukraine cIoffe Physical-Technical Institute, St.Petersburg 194021, Russia

Measurements of FMR in both in-plane (H_{∥}) and out-of-plane (H_{⊥}) magnetic field configurations have been performed on the epitaxial (La_{0.7}Ca_{0.3})_{1−x}Mn_{1+x}O_{3} films. The FMR line in the field H_{∥} is nearly symmetrical from 300 to 4.2 K, while the FMR signal in the perpendicular field has the Dysonian-like shape. The FMR linewidth increases abruptly near the FM transition temperature. Temperature dependences of resonance field H_{res} are well described by the Brillouin function. The angular dependences of H_{res} were found to be in a reasonable agreement with the linear Landau -. Lifshitz equation of motion. By comparing H_{res} for orientations parallel and perpendicular to the film plane, we conclude that the effective anisotropy of the films has the easy-plane character. From the temperature and angular dependences of FMR spectra, the main magnetic parameters were estimated. This work was partly supported by Polish Grant No. 2 PO3B 139 18 and by the Russian Foundation for Basic Research Grant No. 01-02-17479.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : Prof. V.P. Dyakonov

Address for correspondence : ON 3.2 Institute of Physics Polish Academy of Sciences Al. Lotników 32/46 02-668 Warszawa, Poland

Email address : dyakon@ifpan.edu.pl